# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v2

@@ Abstract
Explainable AI (XAI) is a promising route to comply with the EU AI Act, the first multinational AI regulation. XAI enhances transparency and human oversight of AI systems, especially ”black-box“ models criticized as incomprehensible. Yet discourse about the AI Act’s stakeholders and XAI remains disconnected: XAI increasingly prioritizes end users’ needs, while the AI Act focuses on providers’ and deployers’ obligations. We aim to bridge this divide and offer practical guidance on their relationship. Through interdisciplinary discussion in a cross functional team of XAI, AI Act, legal, and requirements-engineering experts, we outline steps to analyze an AI-based clinical decision support system, clarify end-user needs, and assess AI Act applicability. Using an AI system under development as a case study, we show how XAI techniques can help reconcile stakeholder needs with AI Act requirements and fill gaps between usability and regulatory demands. We compare similarities and differences between legal obligations and end-user needs, identify tensions, and point to concrete design choices and trade-offs. We invite researchers and practitioners in XAI to reflect on their role relative to the AI Act and to develop mutual understanding across disciplines. While XAI can help implement core AI Act principles such as transparency and human oversight, it should be considered one element of a broader compliance strategy that also requires standardization, legal interpretation, documentation, organizational processes, governance, testing, and ongoing monitoring and auditing practices. Our findings yield actionable recommendations for integrating XAI into product development, compliance workflows, and stakeholder communication, informing policy-making and standards development.
Keywords European Artificial Intelligence Act, Explainable Artificial Intelligence, Clinical Decision Support, Healthcare, Interdisciplinary Research
@@ 1 Introduction
After long and complicated negotiations between the European Parliament, the EU Commission, and the EU Member States, the AI Act entered into force in August 2024 (aiact). AI-based systems are now regulated under a comprehensive set of product safety rules for the first time. In the coming years, different parts of the EU AI Act will become applicable. Providers and deployers of AI systems need to learn how to apply the different articles of the act, while notified bodies and certification authorities will need to establish a process to assess these systems according to the regulations. Many standardization efforts are already underway, but there will be significant uncertainty about how the AI Act has to be applied and how it relates to different aspects of AI currently under research.
Regulations such as the AI Act are especially important in high-risk artificial intelligence (AI) systems to ensure that they are safe and respect fundamental rights and values (van_kolfschooten_eu_2024). The AI Act follows a risk-based approach: the higher the risk, the stricter the rule. Especially in high-risk systems, it describes many obligations for the provider and the deployer of AI systems.
@@
cdss, which may fall under the definition of a high-risk AI system in the EU AI Act, place the responsibility for the high-risk decision on the end-user (typically, clinicians). However, the AI Act does not directly address end-user needs and instead primarily focuses on regulating the providers and deployers of AI systems. One of the few mentions of end-users is the statement that high-risk AI systems must be designed so that humans can effectively oversee these systems. However, there are no technical requirements that prioritize the user’s needs at the center of its goals (van_kolfschooten_eu_2024).
In contrast, nowadays more explainable artificial intelligence (XAI) researchers suggest putting end-users and their needs at the center of attention (langer_what_2021). XAI methods provide descriptions of the outputs of opaque AI models, thereby satisfying the needs of the end-users. This is very helpful in high-risk applications such as healthcare. Previous work has started exploring the role that XAI plays w.r.t. ensuring the trustworthiness of high-risk applications under the EU AI Act (panigutti_role_2023; kiseleva_transparency_2022; longo_explainable_2024; esmaeilzadeh_challenges_2024; bibalLegalRequirementsExplainability2021). XAI can contribute to implementing transparency and human oversight, which are key principles of the AI Act (panigutti_role_2023). Taking this line of thinking further, we believe that XAI—if designed correctly—could play a crucial role in fulfilling the obligations of deployers and providers of AI systems. However, we acknowledge that XAI alone cannot fulfill all obligations under the AI Act. Full compliance requires additional measures such as the development of harmonized standards, risk management frameworks, and legal assessments, which are beyond the scope of XAI methods. Therefore, we ask:
@@
To answer the research question, we analyze the relationship between the end-user’s XAI needs and the AI Act implications from a legal and technical perspective. We do this by using a concrete, AI-based clinical decision support system (CDSS) under development. This real-world application is currently under development in collaboration with clinicians and academic partners. The CDSS helps intensive care unit (ICU) coordinators who make choices about which patients can be moved out of the ICU with predicted key indicators such as the remaining length of stay (LOS). It is important to note that clinicians can use the predictions to make a decision in our CDSS — the system never decides on its own. To overcome acceptance and trust issues, we use XAI methods to provide human-understandable explanations of the system.
Research on the AI Act and research on XAI are often conducted independently, lacking interdisciplinary collaboration (bexAILawTransdisciplinary2025; gal_bridging_2023). By providing a joint understanding of the role of XAI in the AI Act from an interdisciplinary perspective representing various stakeholder needs, we…
@@
    provide an analysis of the implications of the AI Act in relation to adopting AI in healthcare (see Section˜4)
@@
    describe how the needs of the CDSS stakeholders can be captured through XAI (see Section˜5)
@@
With this, our work is the first that analyzes the role of XAI in the AI Act in a cross-functional team, on a real-world CDSS that takes user-centered needs into account and contributes to the state-of-the-art by mapping obligations of the EU AI Act with the obligations of stakeholders and how to fulfill them using XAI. With this, we encourage researchers and practitioners from the XAI community to reflect on their role in relation to the AI Act by bridging the gap with end-users.
@@ 2 Related Work and Background
In this section, we give an overview of the AI Act and XAI. We also discuss relevant related work. Several studies have already addressed the question of how AI can be applied in healthcare, incorporating ethical and legal considerations (aboy_navigating_2024; porsdam_mann_eu_2024; karimian_ethical_2022; kiseleva_transparency_2022). The function of XAI in fostering transparency and human oversight has also been the subject of considerable debate (panigutti_role_2023). Other works have dealt with the role of XAI in the AI Act (e.g., (gal_bridging_2023)) and have given an overview of the AI Act and XAI (e.g., (panigutti_role_2023)). Nevertheless, it remains unexplored how XAI can be effectively utilized to fulfill the obligations of stakeholders as defined in the AI Act.
@@ 2.1 Background: The EU AI Act
In 2024, the European Union ratified the regulation of AI as technology (aiact). The new regulation was motivated by a need to create trust for AI with respect to product safety, health, and fundamental rights. The AI Act thus incorporates AI systems into the EU’s existing framework for product safety. Certain AI systems will need to undergo conformity assessment and subsequently carry the CE-mark as a sign of being a certified product, just like elevators, machinery, and medical devices (see Section˜2.1).
The regulation also gives authorities an important role in guaranteeing the responsible development and usage of AI, as public sector bodies are to conduct market surveillance but also foster AI innovation (see e.g., (regulatingtrust)). There are two central actors from the perspective of the regulation: the provider (see Section˜2.1) and the deployer (see Section˜2.1) of an AI system. These actors are obliged to take on responsibility for product safety.
CE-marking AI Systems
The conformity assessment for CE-marking AI systems can then be done in two ways. The first is through self-assessment, where the provider of the system certifies that the requirements are met. This is also referred to as internal conformity assessment. The second conformity assessment procedure involves a notified body, an organization that has been accredited to be competent in conformity assessment of certain products. The role of the notified body is to give a neutral assessment of the extent to which the requirements are met. The nature of the AI system will determine which process is to be applied (conformityassessment).
Obligations of the Provider
@@
There will be harmonized standards issued by the European standardization organizations CEN and CENELEC that further detail the requirements in Art. 9 to 15 and 17. An overview of the standardization effort can be found on the two organizations’ joint technical committee’s website (europeancommitteeforstandardizationCENCENCLC2025).
Obligations of the Deployer
@@ 2.2 Background: XAI Terminology
Explainability, Interpretability, and Transparency
The terms explainability, interpretability, and transparency are often used interchangeably (al-ansari_user-centered_2024; renftle_what_2024; hoffman_measures_2023; barredo_arrieta_explainable_2020; mohseni_multidisciplinary_2021). We distinguish these terms by categorizing them into approaches like Explainable AI and Interpretable AI, and desired properties such as Interpretability, Transparency, and Explainability (mohseni_multidisciplinary_2021). Explainable AI provides human-understandable explanations for predictions made by complex AI systems that are not inherently interpretable (wang_designing_2019). Interpretable AI, on the other hand, requires no additional methods for human understanding, as these systems are inherently interpretable (mohseni_multidisciplinary_2021). Interpretability refers to the ability to explain or provide meaning in understandable terms, while Transparency indicates a model is interpretable by design (barredo_arrieta_explainable_2020).
Stakeholder Terminology
XAI goals such as trustworthiness, informativeness, and fairness have been widely studied (al-ansari_user-centered_2024; renftle_what_2024; hoffman_measures_2023; barredo_arrieta_explainable_2020; mohseni_multidisciplinary_2021). Recently, these goals have shifted to focus more on specific stakeholders (samek_towards_2019; biecek2024position), their desiderata (langer_what_2021), and expertise levels (mohseni_multidisciplinary_2021). Stakeholders’ interests, goals, expectations, needs, and demands regarding AI systems are referred to as their desiderata (langer_what_2021).
Understanding different stakeholders in XAI is essential for creating systems that meet various desiderata (langer_what_2021). Different groups have distinct interests, and recognizing these differences is key to developing methods that align with their specific desiderata. The success of an XAI method relies on satisfying stakeholders’ desiderata, which drives explainability approaches that enhance understanding. langer_what_2021, following barredo_arrieta_explainable_2020, categorize stakeholders into five classes: users, (system) developers, affected parties, deployers, and regulators. This classification ensures all relevant needs are considered, with the challenge being to provide suitable explanations for each group (mohseni_multidisciplinary_2021).
Following mohseni_multidisciplinary_2021, end-users are divided into three groups based on their expertise level: Novice Users, who use AI systems daily with little machine learning (ML) knowledge; Data Experts, including data scientists and domain experts using ML for analysis; and AI Experts, comprising ML scientists and engineers designing and implementing ML algorithms and XAI methods.
Taxonomy of Explainability Methods
We adopt the taxonomy proposed by speith_review_2022 as it provides a structured and widely used framework for categorizing XAI methods. These categories are later applied in our methodology (see Section˜5.2) to guide the selection of suitable XAI techniques for our CDSS.
Stage refers to when the XAI method is applied. Ante-hoc methods are inherently interpretable AI models due to their simplicity. Post-hoc methods explain non-transparent ML models, like deep neural networks, using independent algorithms to provide feature relevance or visual explanations.
Applicability refers to post-hoc methods that are either model-specific, using the model’s internal structure, or model-agnostic, applicable to any ML model.
Scope refers to local explanations that detail specific input-output relations, while global explanations describe the overall ML model. Examples include model visualizations (liu_towards_2017) and decision rules (lakkaraju_interpretable_2016).
Other taxonomies include input data (e.g., image, time series), output format (e.g., numerical, visual), and problem type (e.g., classification, regression).
@@ 2.3 Related Work
As previously mentioned, several studies address the question of how AI can be applied in healthcare in general. Table˜1 provides an overview of the related work papers and indicates how we position our own contribution.
Table 1: The related work can be divided into work that explores the role of XAI in the AI Act and work that analyzes the AI Act in healthcare from XAI or user perspectives. Our work extends these preceding studies by combining the user-centric view with the XAI view on the AI Act in a real-world healthcare scenario.
Paper
AI Act
XAI
Healthcare
Real World
User-Centric
gal_bridging_2023
pavlidis_unlocking_2024
fresz_how_2024
onitiu_limits_2023
solaiman_regulating_2024
van_kolfschooten_eu_2024
panigutti_role_2023
@@
gal_bridging_2023 explore the transparency gap between the AI Act and XAI understanding, emphasizing the differing conceptualizations. They focus on stakeholder engagement and how XAI can address these concerns, aligning with our approach. Identified solutions include concentrating on legal definitions, AI system output, and data-related requirements to meet transparency standards. It is important to note that transparency requirements are broad and should not be seen as concrete engineering instructions. XAI contributes to transparency but should not be viewed as a standalone solution; instead, it facilitates transparency while adhering to its foundational principles.
pavlidis_unlocking_2024 explores the role of XAI as a tool to bridge the gap between the opaque nature of ML decisions and the human demand for comprehensibility. In essence, XAI serves as a mechanism to align the objectives of the AI Act, such as enhancing transparency, with the preferences of end-users. The author examines the integration of XAI into the AI Act, with a focus on critical issues like standard setting, oversight, and enforcement. fresz_how_2024 review legal requirements for explainability of AI systems from the AI Act. Furthermore, they describe legal bases for the use of XAI and derive requirements that the XAI methods need to fulfill. All three articles deal with the role of XAI in the AI Act. Our work extends this view by elaborating the role of XAI in a real-world healthcare scenario and by incorporating the end-user needs.
A more domain-specific approach is presented by onitiu_limits_2023, who aims to translate technical safeguards in Art. 13-14 of the AI Act proposal for the use of medical diagnostic systems. The author proposes a broader perspective on the socio-legal implications of technical safeguards, such as post-hoc explainability for decision-making. onitiu_limits_2023 emphasizes that post-hoc XAI methods must focus on harmonizing human expertise with patient-oriented values.
Similarly, solaiman_regulating_2024 explore the AI Act in the healthcare domain with a user-centric focus on AI. Hereby, XAI plays a subordinate role. Instead, the authors focus on the doctor-patient relationship that evolved from a relation where the doctors held the dominant role and the patient follows their guidance without explanation, to a relationship where the patient is seen as an active participant in their healthcare decisions. AI and especially XAI can have an important impact providing more information to both clinicians and patients, thus enhancing patient autonomy.
van_kolfschooten_eu_2024 analyze the implications of the AI Act for the healthcare domain. Here, they give an overview of the obligations for various healthcare stakeholders. Then they use their insights to provide a guideline to address the unique needs of the stakeholders and make recommendations for the implementation and standardization of the act. They also highlight the importance of collaborating with different stakeholders like AI developers, healthcare professionals, patient communities, and regulators to develop those guidelines.
The preceding three studies have all examined the AI Act in the context of healthcare from the perspective of users and XAI. In contrast to these studies, our approach involves a combination of both user-centric and XAI perspectives, which entails the analysis of healthcare stakeholders from the perspectives of AI Act and XAI.
In a more application-oriented study, panigutti_role_2023 conduct an interdisciplinary analysis of the AI Act’s approach to opaque AI systems, aiming to clarify XAI’s role. They focus on Art. 13, concerning transparency, and Art. 14, regarding human oversight, providing an overview of XAI terminology and key concepts. The authors also review the XAI field, noting challenges and limitations consistent with previous findings. They analyze how the AI Act addresses AI system opacity through transparency and human oversight, presenting a use case involving an AI-based proctoring system without XAI tools. In contrast, our study extends beyond Art. 13 and 14 and examines other articles that establish obligations for healthcare stakeholders.

### v3

# The Bidirectional Relationship Between XAI and Regulation: Operationalizing XAI for the AI Act

@@ Abstract
The EU AI Act makes explainability urgent for high-risk AI systems, yet most XAI research focuses on technical metrics rather than regulatory compliance. Understanding how legal requirements reshape XAI method design is challenging: the AI Act regulates organizational relationships (providers, deployers) using legal terminology, specifies obligations without concrete technical requirements, and underrepresents end-users — the very stakeholders whose needs human-centered XAI addresses. As regulations emerge globally, human-centered XAI practitioners face both a challenge and an opportunity: regulations pull XAI research toward real-world deployment, while practitioners can actively shape how explainability enables compliance. This establishes a bidirectional relationship.
Our contribution is threefold. First, we provide the first interdisciplinary analysis of XAI’s role in the AI Act—conducted by a team comprising AI Act legal experts, ML engineers, and requirements engineers — on a real-world clinical decision support system. Second, we systematically align XAI stakeholder roles with AI Act legal responsibilities, revealing where explainability methods address regulatory requirements versus where additional measures are necessary. Third, we identify three key opportunities for human-centered XAI practitioners: actively defining their roles in regulatory implementation; making the user-to-affected-party relationship explicit where regulations address only provider-deployer obligations; and enabling compliance while building multi-level trust — from regulators to affected parties.
Keywords European AI Act $\cdot$ Explainable AI $\cdot$ Human-centered Explainable AI $\cdot$ Regulation $\cdot$ Healthcare $\cdot$ Interdisciplinary Research
Paper accepted at the Fourth World Conference on Explainable Artificial Intelligence, xAI 2026, Fortaleza, Brazil, July 1-3, 2026.
@@ 1 Introduction
After long and complicated negotiations between the European Parliament, the EU Commission, and the EU Member States, the AI Act entered into force in August 2024 (aiact). AI-based systems are now regulated under a comprehensive set of product safety rules for the first time. In the coming years, different parts of the EU AI Act will become applicable. Providers and deployers of AI systems need to learn how to apply the different articles of the act, while notified bodies and certification authorities will need to establish a process to assess these systems according to the regulation. Many standardization efforts are already underway, but there will be significant uncertainty about how the AI Act has to be applied and how it relates to different aspects of AI currently under research.
Regulations such as the AI Act are especially important in high-risk artificial intelligence (AI) systems — such as those in healthcare — to ensure that they are safe and respect fundamental rights and values (Van Kolfschooten and Van Oirschot, 2024). The AI Act follows a risk-based approach: the higher the risk, the stricter the rule. For high-risk systems, it describes many obligations for the provider and the deployer of AI systems.
@@
cdss, which may fall under the definition of a high-risk AI system in the EU AI Act, place the responsibility for the high-risk decision on the end-user (typically, clinicians). However, the AI Act does not directly address end-user needs and instead primarily focuses on regulating the providers and deployers of AI systems. One of the few mentions of end-users is the statement that high-risk AI systems must be designed so that humans can effectively oversee these systems. There are no technical requirements that prioritize the user’s needs at the center of its goals (Van Kolfschooten and Van Oirschot, 2024).
In contrast, nowadays [^1] This has not always been the case. A human-centered perspective on explainable artificial intelligence (XAI) — one that incorporates stakeholder needs, integrates insights from social sciences, and critically reflects on XAI’s capabilities and limitations — has only emerged in recent years. We focus exclusively on this approach (see Section 2.2) more XAI researchers suggest putting end-users and their needs at the center of attention (Langer et al., 2021). XAI methods provide descriptions of the outputs of opaque AI models, thereby satisfying the needs of the end-users. This is very helpful in high-risk applications such as healthcare. Previous work has started exploring the role that XAI plays w.r.t. ensuring the trustworthiness of high-risk applications under the EU AI Act (Panigutti et al., 2023; Kiseleva et al., 2022; Longo et al., 2024; Esmaeilzadeh, 2024; Bibal et al., 2021). XAI can contribute to implementing transparency and human oversight, which are key principles of the AI Act (Panigutti et al., 2023).
This creates both demands and opportunities for the XAI community, establishing a bidirectional relationship: providers and deployers need transparency mechanisms to meet compliance obligations, while XAI practitioners have the tools to actively represent end-user perspectives and shape how legal requirements translate into stakeholder-centered implementations. Taking this line of thinking further, we believe that XAI — if designed correctly — could play a crucial role in fulfilling the obligations of deployers and providers of AI systems. However, we acknowledge that XAI alone cannot fulfill all obligations under the AI Act. Full compliance requires additional measures such as the development of harmonized standards, risk management frameworks, and legal assessments, which are beyond the scope of XAI methods.
Therefore, we ask:
@@
To answer the research question, we analyze the relationship between the end-user’s XAI needs and the AI Act implications from a legal and technical perspective. We do this by using a concrete, AI-based clinical decision support system (CDSS). This real-world application is currently under development in collaboration with clinicians and academic partners. The CDSS helps intensive care unit (ICU) coordinators who make choices about which patients can be moved out of the ICU with predicted key indicators such as the remaining length of stay (LOS). It is important to note that clinicians can use the predictions to make a decision in our CDSS — the system never decides on its own. To overcome acceptance and trust issues, we use XAI methods to provide human-understandable explanations of the system.
Research on the AI Act and research on XAI are often conducted independently, lacking interdisciplinary collaboration (Bex, 2025; Gyevnar et al., 2023). By providing a joint understanding of the role of XAI in the AI Act from an interdisciplinary perspective representing various stakeholder needs, we…
@@
    provide an analysis of the implications of the AI Act in relation to adopting AI in healthcare (see Section˜4);
@@
    describe how the needs of the CDSS stakeholders can be captured through XAI (see Section˜5); and
@@
Our work is the first that analyzes the role of XAI in the AI Act in an interdisciplinary team, on a real-world CDSS that takes user-centered needs into account and contributes to the state-of-the-art by mapping obligations of the EU AI Act with the obligations of stakeholders and how to fulfill them using XAI. Through this analysis, we demonstrate how the XAI community can move from passive observers to active participants in regulatory implementation — positioning human-centered explainability as essential to compliance while ensuring that regulations pull XAI research toward real-world stakeholder needs.
@@ 2 Related Work and Background
In this section, we give an overview of the AI Act and XAI. We also discuss relevant related work.
@@ 2.1 Background: The EU AI Act
In 2024, the European Union ratified the regulation of AI as technology (aiact). The new regulation was motivated by a need to create trust for AI with respect to product safety, health, and fundamental rights. The AI Act incorporates AI systems into the EU’s existing framework for product safety. Certain AI systems require conformity assessment and subsequently carry the CE-mark as a sign of being a certified product, just like elevators, machinery, and medical devices.
The regulation also gives authorities an important role in guaranteeing the responsible development and usage of AI, as public sector bodies are to conduct market surveillance but also foster AI innovation (see e.g., regulatingtrust). There are two central actors from the perspective of the regulation: the provider and the deployer of an AI system. These actors are obliged to take on responsibility for product safety.
CE-marking AI Systems.
The conformity assessment for CE-marking AI systems can be done in two ways. The first is through self-assessment, where the provider of the system certifies that the requirements are met. This is also referred to as internal conformity assessment. The second conformity assessment procedure involves a notified body, an organization that has been accredited to be competent in conformity assessment of certain products. The role of the notified body is to give a neutral assessment of the extent to which the requirements are met. The nature of the AI system will determine which process is to be applied (conformityassessment).
Obligations of the Provider.
@@
There will be harmonized standards issued by the European standardization organizations CEN and CENELEC that further detail the requirements in Art. 9 to 15 and 17. An overview of the standardization effort can be found on the two organizations’ joint technical committee’s website (European Committee for Standardization, 2025).
Obligations of the Deployer.
@@ 2.2 Background: Explainable AI
Scope of XAI in This Work.
Broad XAI goals such as trustworthiness, informativeness, and fairness have been widely studied (Amparore et al., 2021; Deck et al., 2024b, a; Hoffman et al., 2023; Barredo Arrieta et al., 2020), but the area of study have recently shifted to focus more on individual stakeholders (Uth et al., 2026; Longo et al., 2024; Biecek and Samek, 2024), their desiderata (Langer et al., 2021), and expertise levels (Mohseni et al., 2021). We explicitly focus on this human-centered XAI — designed with end-user comprehension and stakeholder desiderata in mind. When we refer to “XAI” or “the XAI community” throughout this work, we address researchers and practitioners working within this paradigm. This scoping is essential because only human-centered XAI can meaningfully bridge stakeholder needs and AI Act compliance requirements.
Explainability and Transparency.
The terms explainability, interpretability, and transparency are often used interchangeably in XAI literature (Al-Ansari et al., 2024; Renftle et al., 2024; Hoffman et al., 2023; Barredo Arrieta et al., 2020). The AI Act uses transparency broadly to encompass technical documentation, system capabilities, and information enabling deployers to interpret the system’s output and use it appropriately (Art. 13). We distinguish between explainability as a desired property — the degree to which a human can understand the reasons behind a system’s decisions (Barredo Arrieta et al., 2020) — and XAI as methods that provide human-understandable explanations for predictions made by complex AI systems (Wang et al., 2019; Barredo Arrieta et al., 2020).
Stakeholder Terminology.
XAI success relies on satisfying diverse stakeholder desiderata (Langer et al., 2021; Fairhurst et al., 2026). Langer et al. (2021) categorize stakeholders into five classes: users, developers, affected parties, deployers, and regulators. Mohseni et al. (2021) further divide end-users by expertise: Novice Users (daily AI users with little ML knowledge), Data Experts (data scientists and domain experts using ML for analysis), and AI Experts (ML scientists and engineers implementing ML algorithms and XAI methods).
XAI Method Taxonomy.
We adopt Speith (2022) taxonomy to guide XAI selection for our CDSS (see Section˜5.2). Stage: ante-hoc methods (inherently interpretable models) versus post-hoc methods (explain opaque models with independent algorithms). Applicability: model-specific (using internal structure) versus model-agnostic (applicable to any model). Scope: local explanations (specific input-output relations) versus global explanations (overall model behavior). Other dimensions include input data type, output format, and problem type.
@@ 2.3 Related Work
Several studies have already addressed the question of how AI can be applied in healthcare, incorporating ethical and legal considerations (Aboy et al., 2024; Porsdam Mann et al., 2024; Karimian et al., 2022; Kiseleva et al., 2022). The function of XAI in fostering transparency and human oversight has also been the subject of considerable debate (Panigutti et al., 2023). Other works have dealt with the role of XAI in the AI Act (e.g., (Gyevnar et al., 2023; Juliussen, 2025; Metikoš and Ausloos, 2025)) or the GDPR (e.g., (Juliussen, 2025; Metikoš and Ausloos, 2025)). Nevertheless, how XAI can be effectively utilized to fulfill the obligations of stakeholders as defined in the AI Act remains unexplored. Table˜1 provides an overview of the related work and indicates how we position our own contribution.
Table 1: Related work can be divided into work that explores the role of XAI in the AI Act and work that analyzes the AI Act in healthcare from XAI or user perspectives. Our work extends these studies by combining the user-centric view with the XAI view on the AI Act in a real-world healthcare scenario.
Paper AI Act XAI Healthcare Real World User-Centric
Juliussen (2025)
Buttaboni and Floridi (2026)
Gyevnar et al. (2023)
Pavlidis (2024)
Fresz et al. (2024)
Onitiu (2023)
Solaiman and Malik (2024)
Van Kolfschooten and Van Oirschot (2024)
Metikoš and Ausloos (2025)
Panigutti et al. (2023)
@@
Several studies explore the role of XAI in the AI Act from conceptual and legal perspectives. Juliussen (Juliussen, 2025) provides a comprehensive analysis of the right to explanation under both GDPR and the AI Act, examining XAI methods (intrinsic and post-hoc) in the context of automated decision-making. Buttaboni and Floridi (Buttaboni and Floridi, 2026) propose a regulatory taxonomy that distinguishes transparency, traceability, interpretability, and explainability as layered and interdependent dimensions of AI opacity, providing a conceptual framework for legal interpretation and compliance. Gyevnar et al. (2023) explore the transparency gap between the AI Act and XAI understanding, emphasizing that transparency requirements are broad and should not be seen as concrete engineering instructions; XAI contributes to transparency but should not be viewed as a standalone solution. Pavlidis (2024) examines XAI as a mechanism to align AI Act objectives with end-user preferences, focusing on standard setting, oversight, and enforcement. Fresz et al. (2024) review legal requirements for explainability and derive requirements that XAI methods need to fulfill. Our work extends these studies by elaborating the role of XAI in a real-world healthcare scenario and incorporating end-user needs.
Within healthcare, Onitiu (Onitiu, 2023) translates technical safeguards in Art. 13-14 for medical diagnostic systems, emphasizing that post-hoc XAI methods must harmonize human expertise with patient-oriented values. Solaiman and Malik (Solaiman and Malik, 2024) explore the AI Act in healthcare with a user-centric focus, examining how the doctor-patient relationship has evolved to one where patients are active participants in their healthcare decisions, and how AI and XAI can enhance patient autonomy by providing more information to both clinicians and patients. Van Kolfschooten and Van Oirschot (Van Kolfschooten and Van Oirschot, 2024) analyze AI Act implications for healthcare stakeholders, providing implementation guidelines and highlighting the importance of collaboration with AI developers, healthcare professionals, patient communities, and regulators. In contrast to these studies, our approach combines user-centric and XAI perspectives through an analysis of healthcare stakeholders from both AI Act and XAI viewpoints in a real-world scenario.
Real-world applications are examined by Metikoš and Ausloos (Metikoš and Ausloos, 2025), who analyze GDPR case law on the right to explanation across European judicial bodies and data protection authorities, providing insights into how the AI Act’s right to explanation should be understood in practice. Panigutti et al. (2023) conduct an interdisciplinary analysis focusing on Art. 13 and 14, presenting a use case involving an AI-based proctoring system without XAI tools. In contrast, our study extends beyond Art. 13 and 14 to examine other articles that establish obligations for healthcare stakeholders in a real-world healthcare scenario with integrated user needs.